Suppose a company produces $5M worth of output and has sales of $2M each to domestic and foreign customers. It imports $1 M worth of raw material, pays its workers $3M in wages, pays its creditors $2M in interest, and has minus $1M in profits for its owners. This company’s operations add $4M to GDP whether measured by the value added approach, the expenditure approach, or the income approach.
Explain in detail: True False or Uncertain
1
Expert's answer
2015-02-18T10:13:06-0500
Produces $5M, sales of $2M each to domestic and foreign customers. Imports $1M, $3M in wages, $2M in interest, minus $1M in profits. If
measured by the value added approach the company adds 5M - 1M = $4M, if
measured by the expenditure approach it adds 1M + 2M + 1M = $4M, and if
measured by the the income approach, it adds 3M - 1M + 2M = $4M.
So, this company’s operations add $4M to GDP.
That's why the statement is True.